# Optimal Control to Minimize Dissipation and Fluctuations in Open Quantum Systems Beyond Slow and Rapid Regimes

### v1

@@ IV Conclusion
$W_{\mathrm{diss}}$ $=\beta^{-1}\,S\!\big(\pi(0)\,\|\,\pi(u_{T})\big)\;+\;\int_{0}^{T}\!dt\,\big[u_{T}-u(t)\big]\;R\!\big(u(t)\big),$ (41)
@@
$B(u)$ = $=-\,2\,\gamma(u)\,\coth(\beta u),$ (54)
@@
Appendix B GRAPE algorithm
@@
$p(T)$ $=-\left.\left(\frac{\partial d(T,q)}{\partial q}\right)^{\top}\right|_{q=q(T)},$ (62)
@@
Appendix C calculation of $\partial_{u}A(u)$
@@
Appendix D time evolution of $Y(t)$
In this section we derive the time evolution of the auxiliary operator defined in Eq. (9). By definition (7) and the Leibniz integral rule,
$\partial_{t_{1}}\overleftarrow{\mathcal{P}}(t_{1},t_{2})$ $=\mathcal{L}_{t_{1}}\,\overleftarrow{\mathcal{P}}(t_{1},t_{2}),\qquad\overleftarrow{\mathcal{P}}(t_{2},t_{2})=\mathcal{I}.$ (77)
$\frac{d}{dt}Y(t)$
$=S_{\rho(t)}(\dot{H}(t))+\mathcal{L}_{t}\left[Y(t)\right]$
Then we obtain Eq. (10).
Appendix E Equivalence between the Fock-space description of a single-level quantum dot and a two-level (pseudospin) representation
@@
we may rewrite the projector $|\ket{1}\!\bra{1}$ in terms of $I$ and $\sigma_{z}$:
@@
The term proportional to $I$ produces only an overall energy shift and can be dropped since it does not affect the density-matrix dynamics ($I,\rho]=0$). Defining the rescaled control amplitude

### v2

@@ Appendix A Methods of minimization of fluctuation and dissipation
$W_{\mathrm{diss}}$ $=\beta^{-1}\,S\big(\pi(0)\,\|\,\pi(u_{T})\big)$ (43)
$+\int_{0}^{T}dt\,\big[u_{T}-u(t)\big]R\big(u(t)\big),$
@@
$B(u)$ $=-2\gamma(u)\coth(\beta u),$ (54)
@@
Appendix D GRAPE algorithm
@@
$p(T)$ $=\left.\left(\frac{\partial d(T,q)}{\partial q}\right)^{\top}\right|_{q=q(T)},$ (62)
@@
Appendix E calculation of $\partial_{u}A(u)$
@@
Appendix F Equivalence between the Fock-space description of a single-level quantum dot and a two-level (pseudospin) representation
@@
we may rewrite the projector $|\ket{1}\bra{1}$ in terms of II and $\sigma_{z}$:
@@
The term proportional to II produces only an overall energy shift and can be dropped since it does not affect the density-matrix dynamics ($I,\rho]=0$). Defining the rescaled control amplitude